Section 01  Logic

Section 01: Logic

Invited Lectures

Tuesday, 25 August  H 1012

16:15 – 17:00 Matthew Foreman, University of California, Irvine, USA Generic Large Cardinals: New Axioms for Mathematics? (I: p.30) Chair: Peter G. Koepke

17:15 – 18:00 Ludomir Newelski, University of Wrocław, Wrocław and Polish Academy of Sciences, Poland Meager forking and m-independence (I: p.33) Chair: Alex James Wilkie

Wednesday, 26 August  MA 005

14:00 – 14:45 Stevo B. Todorcevic, CNRS, Université de Paris 7, Paris, France, Mathematicki Institut, Beograd, Yugoslavia, and University of Toronto, Toronto, Canada Basis Problems in Combinatorial Set Theory (I: p.35) Chair: Martin Goldstern

Wednesday, 26 August  H 1012

16:15 – 17:00 Gregory Hjorth, UCLA, Los Angeles, USA When is an equivalence relation classifiable? (I: p.32) Chair: Ronald B. Jensen

17:15 – 18:00 Alex James Wilkie, University of Oxford, Oxford, England O-minimality (I: p.35) Chair: Anand Pillay

Short Communications and Poster Sessions

Short Communications  Thursday, 20 August
Lattices and Universal Algebra  MA 042

Chair: Marcus A. Kracht

14:00 – 14:15 Isabella C. Burger, Rand Afrikaans University, Auckland Park (Johannesburg), South Africa Boolean algebras and information-dependent entailment relations (p.6)

14:15 – 14:30 Dietmar Dorninger, Technische Universität Wien, Wien, Austria On ring-like structures occurring in axiomatic quantum mechanics (p.7)

14:30 – 14:45 Philippe Niederkorn, Université de Liège, Liège, Belgium Natural dualities for finitely generated varieties of MV-algebras (p.8)

Logic  Section 01

Short Communications  Friday, 21 August
Categories: Logic and Topology  MA 042

Chair: Gerhard Preuß

14:00 – 14:15 Michel Hebert, The American University in Cairo, Cairo, Egypt Syntactic view on locally generated categories: abstract and concrete cases (p.7)

14:15 – 14:30 Dieter Leseberg, TU Braunschweig, Braunschweig, Germany Filterconvergence as a convenient concept for Topology (p.8)

Set Theoretic Topology  MA 042

Chair: Dieter Leseberg

15:00 – 15:15 Gerhard Preuß, Freie Universität, Berlin, Germany Convenient Topology — A New Branch of Topology (p.9)

15:15 – 15:30 Harald Nusser, Freie Universität, Berlin, Germany Completions of probabilistic semiuniform convergence spaces (p.8)

Short Communications  Monday, 24 August
Constructive Mathematics and Models of Arithmetic  MA 042

Chair: Hiloshi Ohno

16:15 – 16:30 Douglas S. Bridges, University of Waikato, Hamilton, New Zealand Ideals in constructive ring and Banach algebra theory (p.6)

16:30 – 16:45 Luminita Simona Dediu, University of Waikato, Hamilton, New Zealand A constructive extension theorem for ultraweakly continuous linear functionals (p.6)

16:45 – 17:00 Hiloshi Ohno, Kashiwa, Japan A certain proposition unprovable in number-theoretic formal system (p.9)

Recursion Theory and Model Theory  MA 042

Chair: Helmut Pfeiffer

17:15 – 17:30 David Hartz, College of St. Benedict, St. Joseph, MN, USA Hierarchy Based on Inductive Definitions (p.7)

17:30 – 17:45 Cyrus F. Nourani, METAAI and UCSB, Cardiff By The Sea, CA, USA Functorial Computability And Generic Definable Models (p.8)

Short Communications  Tuesday, 25 August
Set Theory and Recursion Theory  MA 042

Chair: Robert Fittler

14:00 – 14:15 Philip D. Welch, Kobe University, Kobe, Japan The Length of Infinite Time Turing Machine Computations (p.11)

14:15 – 14:30 Martin Zeman, Math. Inst., Humboldt–Universität, Berlin, Germany ◇ at Mahlo Cardinals (p.11)

14:30 – 14:45 Marat-M. Arslanov, Kazan-State-University, Kazan, Russia The enumeration degree structures (p.5)

Model Theory and Recursion Theory  MA 042

Chair: Andreas Baudisch

15:00 – 15:15 Mikhail Peretyat’kin, Institute of Pure and Applied Mathematics, Almaty, Kazakhstan Expressive possibilities of finitely axiomatizable theories (p.9)

15:15 – 15:30 Niandong Shi, Dept. of Mathematics, East Stroudsburg, PA, USA Universal graph problem and model theory (p.10)

15:30 – 15:45 Helmut Schwichtenberg, Universität München, München, Germany Ramification, Modality, and Linearity in Higher Type Recursion (p.10)

2  ICM’98 – Program

Section 01  Logic

Poster Session  Monday, 24 August 14:00 – 15:45  MA 144

Fariborz Azarpanah, University of Chamran, Ahvaz, Iran

Sum and intersection of summand ideals in C(X) (p.5) Grzegorz Bancerek, University of Białystok, Białystok, Poland

Mizar — a work environment for a mathematician: an experiment with continuous lattices (p.6) Peter Michael Schuster, Ludwig-Maximilians-Universität München, München, Germany

A Sequence Model for Constructive Nonstandard Analysis (p.10)

ICM’98 – Program  3